How much 5% hcl must be mixed with 60ml of 20% hcl to get a 10% solution?
I think that it would just be another 60ml, but I'm not sure.

Answer by mananth(16946)

--------percent ---------------- quantity
Hcl I 20 ---------------- 60 ml
Hcl II 5 ---------------- x ml
Mixture 10 ---------------- 60 + x ml

20 * 60 + 5 x = 10 ( 60 + x )
1200 + 5 x = 600 + 10 x
5 x + -10 x = 600 - -1200
-5 x = -600
/ -5
x = 120
120 ml of 5 % Hcl II